# Decoupling Strategy and Execution in Task-Focused Dialogue via Goal-Oriented Preference Optimization

### v1

@@ 4.2.1 Advantages of GOPO over Mainstream Training Frameworks
Across all four datasets, GOPO consistently outperforms the untrained model and strong baselines including SFT, PPO, and Memento. Specifically, GOPO improves TSE by 25%–38% and Reward by 12%–40% over the untrained model, demonstrating the effectiveness of goal-oriented training.
Compared to SFT, GOPO achieves 10%–15% higher TSE, highlighting its advantage in sequence-conditioned, business-driven dialogue tasks beyond token-level NLG optimization. Relative to PPO, GOPO yields 7%–10% gains in TSE and 3%–20% in Reward, validating the benefit of dual-agent decomposition and business-aligned reward design over single-agent RL. Finally, GOPO outperforms Memento by 9%–12% in TSE and 3%–8% in Reward, indicating strong robustness across training paradigms.
@@ 4.2.3 Performance Analysis under Different GOPO Parameter Settings
The training trends and experimental results of the GOPO framework under different parameter settings are shown in Figure 3 and Table 4. It can be observed that as the model scale increases, GOPO achieves improvements in TSE, Reward, and BLEU across all datasets, with GOPO-Qwen3-14B performing the best. The medium-sized GOPO-Qwen3-8B also significantly outperforms GOPO-Qwen-7B-Chat, while the performance gains vary across datasets, with more noticeable improvements on Mgshop and the TmallBrand series. Notably, despite variations in individual loss values, the overall reward converges reliably.
@@ 4.3 Ablation Study
Removing the Expert Agent causes substantial performance degradation, confirming its critical role in long-horizon optimization. On TmallBrand-A, TSE drops by 5.7% and Reward decreases by 13.3%. Notably, although BLEU slightly increases, TSE declines sharply, indicating that without strategic guidance the model produces conservative yet inefficient responses. This further exposes the limitation of BLEU as a surface-level metric for task-oriented dialogue.
Compared to the untrained baseline, Full GOPO achieves significant gains, improving TSE by 28.9% and 33.6%, and Reward by 10.3% and 16.0% on the two datasets. These results demonstrate that the ESNDCG-driven hierarchical RL objective effectively guides the model toward efficient task completion.
@@ 5 Conclusion
We address the “efficiency trap” in task-oriented dialogue systems for complex business scenarios, where optimization objectives misalign with business KPIs and strategies lack flexibility and compliance. We propose GOPO, a hierarchical RL framework with a dual-agent architecture that decouples strategy planning from response execution. An Expert Agent selects high-level strategies, while a Customer Service Agent generates compliant responses under strong constraints. Combined with SOP-matching and a joint reward model aligned with business end-state metrics, GOPO shifts optimization from preference imitation to direct business value maximization. Experiments on real-world datasets show consistent gains over all baselines, including large proprietary models, on metrics such as TSE and Reward.
GOPO currently relies on a manually defined policy pool, and its reward design can be further improved with advanced credit assignment. Future work will explore automatic strategy discovery and broader domain generalization. Overall, GOPO offers a practical framework for business-oriented task dialogue and opens new directions for goal-driven multi-agent LLM optimization.

### v2

@@ 4.2.1 Advantages of GOPO over Mainstream Training Frameworks
Model Mgshop TmallBrand-A
TSE GRE BLEU TSE GRE BLEU
Full GOPO 0.9475 7.63 0.279 0.8511 7.46 0.186
GOPO w/o Expert Agent 0.9371 7.16 0.217 0.8023 6.47 0.227
Full Baseline 0.7353 6.92 0.251 0.6371 6.43 0.213
Baseline w/o Expert Agent 0.4327 6.08 0.240 0.1818 5.19 0.182
Table 6: Performance of Customer Service Agents Evaluated with Trained GOPO Expert Agent on Mgshop and TmallBrand-A Datasets.
Customer Service Agent Mgshop Dataset TmallBrand-A Dataset
TSE GRE BLEU TSE GRE BLEU
GOPO-Qwen3-14B 0.9475 7.63 0.279 0.8511 7.46 0.186
Qwen-235B 0.9352 7.45 0.214 0.8428 7.37 0.201
DeepSeek-R1 0.9453 7.55 0.242 0.8503 7.03 0.245
Gemini-2.5 0.9268 7.52 0.220 0.8025 6.83 0.131
Table 7: Performance of Fixed GOPO Customer Service Agent with Different Expert Agents on Mgshop and TmallBrand-A Datasets.
Expert Agent Mgshop Dataset TmallBrand-A Dataset
TSE GRE BLEU TSE GRE BLEU
GOPO-Qwen3-14B 0.9475 7.63 0.279 0.8511 7.46 0.186
Qwen-235B 0.8714 7.40 0.193 0.8291 6.27 0.203
DeepSeek-R1 0.7854 6.73 0.234 0.8395 6.84 0.146
Gemini-2.5 0.7384 5.30 0.251 0.8163 6.55 0.228
Across all four datasets, GOPO consistently outperforms the untrained model and strong baselines including SFT, PPO, and Memento. Specifically, GOPO improves TSE by 23%–26% and GRE by 23%–40% over the untrained model, demonstrating the effectiveness of goal-oriented training.
Compared to SFT, GOPO achieves 7.5%–10% higher TSE, highlighting its advantage in sequence-conditioned, business-driven dialogue tasks beyond token-level NLG optimization. Relative to PPO, GOPO yields 0.6%–7.7% gains in TSE and 3%–20% in GRE, validating the benefit of dual-agent decomposition and business-aligned GRE design over single-agent RL. Finally, GOPO outperforms Memento by 9%–12% in TSE and 3%–8% in GRE, indicating strong robustness across training paradigms.
@@ 4.2.3 Performance Analysis under Different GOPO Parameter Settings
The training trends and experimental results of the GOPO framework under different parameter settings are shown in Figure 3 and Table 4. It can be observed that as the model scale increases, GOPO achieves improvements in TSE, GRE, and BLEU across all datasets, with GOPO-Qwen3-14B performing the best. The medium-sized GOPO-Qwen3-8B also significantly outperforms GOPO-Qwen-7B-Chat, while the performance gains vary across datasets, with more noticeable improvements on Mgshop and the TmallBrand series. It is worth noting that the loss value of the Expert Agent remains relatively high and decreases only marginally during training. This behavior is expected given that the Expert Agent is optimized with a policy-gradient objective augmented by entropy regularization, where the loss magnitude reflects both non-trivial advantage estimates and sustained policy stochasticity rather than convergence to zero. Notably, despite the Expert loss stabilizing at a higher value, the overall reward converges reliably across all settings, indicating effective long-horizon optimization.
@@ 4.3 Ablation Study
Removing the Expert Agent causes substantial performance degradation, confirming its critical role in long-horizon optimization. On TmallBrand-A, TSE drops by 5.7% and GRE decreases by 13.3%. Notably, although BLEU slightly increases, TSE declines sharply, indicating that without strategic guidance the model produces conservative yet inefficient responses. This further exposes the limitation of BLEU as a surface-level metric for task-oriented dialogue.
Compared to the untrained baseline, Full GOPO achieves significant gains, improving TSE by 28.9% and 33.6%, and GRE by 10.3% and 16.0% on the two datasets. These results demonstrate that the ESNDCG-driven hierarchical RL objective effectively guides the model toward efficient task completion.
@@ 5 Conclusion
This paper addresses misalignment in task-oriented dialogue systems under complex business scenarios. We propose GOPO, a hierarchical RL framework with dual agents, separating strategy planning from response generation. The Expert Agent chooses sequences of skills, while the Customer Service Agent executes responses following strict compliance, guiding the system toward desired business outcomes. GOPO combines SOP compliance with a joint reward over entire dialogues, shifting optimization from preference imitation to direct business value maximization. Experiments on multiple real-world datasets show GOPO consistently outperforms strong baselines, including large proprietary models on TSE and GRE metrics. Current evaluation partly relies on offline TSE from real dialogues, with ongoing online tests confirming effectiveness and stability. Future work includes automated policy and skill discovery, expanding the skill pool dynamically, and adapting across domains. Overall, GOPO provides a practical and extensible framework for task-oriented dialogue systems and opens new directions for optimizing large multi-agent models toward specific objectives.